# Quantum-inspired activation functions in the convolutional neural network

### v1

@@ Abstract
Driven by the significant advantages offered by quantum computing, research in quantum machine learning has increased in recent years. While quantum speed-up has been demonstrated in some applications of quantum machine learning, a comprehensive understanding of its underlying mechanisms for improved performance remains elusive. Our study fills this gap by examining the expressibility of quantum circuits integrated within a convolutional neural network (CNN). Through numerical training on the MNIST dataset, our hybrid quantum-classical CNN model exhibited superior feature selection capabilities and significantly reduced the required training steps compared to the classical CNN. To understand the root of this enhanced performance, we conducted an analytical investigation of the functional expressibility of quantum circuits and derived a quantum activation function. We demonstrated that this quantum activation is more efficient in selecting important features and discarding unimportant information of input images. These findings not only deepen our comprehension of quantum-enhanced machine-learning models but also advance the classical machine-learning technique by introducing the quantum-inspired activation function.
Introduction
The unparalleled potential of quantum algorithms over their classical counterparts has ignited widespread enthusiasm for quantum computing [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Thanks to the achievements in quantum hardware development [11, 12, 13, 14, 15], quantum supremacy was demonstrated via the random sampling task made by the Google superconducting quantum computer and the photonic quantum device [16, 17]. In addition to these landmark achievements, quantum computing has found applications across diverse domains, including simulating many-body Hamiltonian [18, 19, 20, 21, 2, 22, 23, 24], simulating spectroscopies [25, 26, 27, 28, 29], and solving NP-complete problems [30, 31, 32, 33, 34]. Among them, most of these applications use the variational technique to find a solution with a minimal value of the loss function or the energy, which is analogous to the strategy used in machine learning techniques, inspiring the idea of quantum machine learning [35, 36].
Figure 1: Five quantum circuits used in our study. The orange- and green-colored rectangles represent parametric $R_{y}$ rotation operators. In circuits 2, 4, and 5, these two $R_{y}$ operators are separated by a CNOT gate. In circuit 1 and circuit 3, two $R_{y}$ operators are applied one after another, which can be merged into a single $R_{y}$ operator.
@@ Introduction
Understanding the origin of superior performance in quantum machine learning is essential for optimizing quantum models and potentially enriching classical algorithms. Quantum-inspired classical algorithms, for example, have already accelerated linear algebra computations, showcasing the benefits of translating quantum insights to classical contexts [58, 59, 60, 61]. Despite these advancements, the machine learning field sees a paucity of quantum-inspired algorithms, primarily due to an inadequate grasp of quantum model intricacies [62]. This knowledge gap has spurred investigations into the expressibility of quantum models [63, 64, 65, 66]. However, giving an analytical relationship between the expressibility and the architecture of a quantum circuit is challenging, and this difficulty increases with the number of qubits and quantum gates. Consequently, research has predominantly pursued numerical evaluations, including evaluating the expressibility of the Hilbert space and the functional expressibility [63, 64, 65, 66]. For instance, the expressibility of the Hilbert space was characterized by comparing different distributions of the fidelity between the quantum state from a quantum model and the Haar-random states [63]. In addition, it was proposed to understand the functional expressibility of a quantum model by utilizing the Fourier series [64]. While these insights advance our understanding, they fail to elucidate the relationship between Hilbert space expressibility and quantum machine learning efficacy. Moreover, the decomposition of a function into a Fourier series cannot intuitively reflect the nonlinear relationship between the input data and the output. To advance quantum machine learning techniques, it is vital to address these foundational questions.
Figure 2: Hybrid quantum-classical convolutional neural network (CNN). a, A figure for the hybrid CNN architecture, which includes one convolution layer, one average pooling layer, one flattened layer, and two fully connected (FC) neural networks. b, A quantum circuit with nine qubits is sketched. c, A figure for advancing the classical machine learning method via quantum-inspired algorithms.
Figure 3: The performance of the hybrid convolutional neural network (CNN) and the classical CNN. The top panel plots the loss function at each epoch during the training phase. The bottom panel plots the training accuracy. The MNIST dataset with 57,600 samples is used to train these CNN models.
Our research advances the understanding of quantum models in quantum machine learning by providing clear, analytical expressions for quantum circuits, setting our work apart from prior studies. Specifically, we analyze five shallow quantum circuits integrated within a convolutional neural network (CNN) framework. To compare the hybrid convolutional neural network (CNN) and the classical CNN, we restrict ourselves to the situation that these five quantum models have the same number of variables as that in the classical CNN. To understand the trainability of these circuits, we first study the expressibility of the Hilbert space of these circuits and observe a similar expressibility and trainability among these five circuits. We then explore the training speed of hybrid CNNs with these five quantum circuits and attribute different performances to different functional expressibilities. A notable discovery is that circuits employing a rotation operator directly after the data upload via angle encoding outperform those incorporating a CNOT layer between the data encoding and rotation operations. This insight is instrumental for optimizing quantum circuit design in hybrid systems. Further, we succeeded in deriving analytical expressions of the quantum circuit yielding the best performance in the hybrid CNN, which is characterized as a continued product of the cosine series. Compared to the traditional Tanh and Sigmoid activation functions, this quantum activation function can efficiently select essential features and discard unimportant information, resulting in faster loss function convergence. This finding is pivotal because we offer a clear explanation for the superior performance of hybrid CNNs and a quantum-inspired activation function with enhanced efficiency to the broader machine-learning community.
Results
Quantum Circuits
Our study aims to elucidate the performance of hybrid quantum-classical machine learning frameworks with different quantum models. To avoid the quantum barren plateau [67], we focus on five shallow quantum circuits (QCs), as depicted in Fig. 1. In QCs 1 and 3, sequential parametric $R_{y}$ rotation operators—represented as orange and green rectangles—proceed without intermediary quantum operations. Conversely, the two rotation operators are separated by a CNOT layer in QCs 2, 4, and 5 (QC 5 is investigated in Ref. 63). These five circuits have different architectures, generating different expressibilities of the Hilbert space and functional expressibilities. As discussed in early studies [63, 65], these two expressibilities could affect the performance of the hybrid quantum-classical neural network, although the hidden mechanism is unclear. To explore this relationship, we will integrate these quantum models into a CNN.
Before performing a machine learning task, we need to understand the trainability of these quantum circuits, which can be characterized by the expressibility of the Hilbert space. The relationship between the trainability and the expressibility was explained in Ref. 68, which concluded that a quantum circuit with strong expressibility of the Hilbert space exhibits the quantum barren plateau issue, leading to a weak trainability. To quantize the expressibility of the Hilbert space of a quantum circuit, we follow Ref. 63 and compare the distribution of quantum states obtained from sampling the parameters of a quantum circuit to the uniform distribution of Haar-random states. The Kullback-Leibler (KL) divergence is used to characterize different distributions of the fidelity of these quantum states (see details in the method section). A smaller value of KL divergence denotes a better expressibility of the Hilbert space. It is found that these five circuits generate a similar distribution of fidelities, leading to a comparable value of KL divergence, which is about 0.345 (see Supplementary Note 1). This result implies that these five quantum circuits have a similarly limited expressibility of the Hilbert space; as a result, they exhibit good trainability [68].
Figure 4: Analysis of Quantum Circuits. a, The quantum state $|\Psi_{1}^{N}\rangle$ of quantum circuit 1 is divided into two parts, $|\Psi_{1}^{2}\rangle$ and $|\Psi_{3}^{N}\rangle$, where $N$ is the total number of qubits. b, The contribution of a $2\times 2$ filter encoded into quantum circuit 1 comes from the first and fourth qubits. c, The contribution of a $3\times 3$ filter encoded into quantum circuit 1 comes from qubits with odd indexes. d, The analysis of quantum circuit 3. e, The contribution of a $2\times 2$ filter encoded into quantum circuit 3 comes from the first and third qubits. f, The contribution of a $3\times 3$ filter encoded into quantum circuit 3 comes from the first, third, and seventh qubits.
Hybrid quantum-classical convolutional neural network

### v2

@@ Abstract
Driven by the significant advantages offered by quantum computing, research in quantum machine learning has increased in recent years. While quantum speed-up has been demonstrated in some applications of quantum machine learning, a comprehensive understanding of its underlying mechanisms for improved performance remains elusive. Our study address this problem by investigating the functional expressibility of quantum circuits integrated within a convolutional neural network (CNN). Through numerical experiments on the MNIST, Fashion MNIST, and Letter datasets, our hybrid quantum-classical CNN model demonstrates superior feature selection capabilities and substantially reduces the required training steps compared to classical CNNs. Notably, we observe similar performance improvements when incorporating three other quantum-inspired activation functions in classical neural networks, indicating the benefits of adopting quantum-inspired activation functions. Additionally, we developed a hybrid quantum Chebyshev-polynomial network (QCPN) based on the properties of quantum activation functions. We demonstrate that a three-layer QCPN can approximate any continuous function, a feat not achievable by a standard three-layer classical neural network. Our findings suggest that quantum-inspired activation functions can reduce model depth while maintaining high learning capability, making them a promising approach for optimizing large-scale machine-learning models. We also outline future research directions for leveraging quantum advantages in machine learning, aiming to unlock further potential in this rapidly evolving field.
I Introduction
The unparalleled potential of quantum algorithms over their classical counterparts has ignited widespread enthusiasm for quantum computing [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Due to achievements in quantum hardware development [11, 12, 13, 14, 15], quantum supremacy was demonstrated via the random sampling task made by Google superconducting quantum computers and photonic quantum devices [16, 17]. In addition to these landmark achievements, quantum computing has found applications across diverse domains, including simulating many-body Hamiltonian [18, 19, 20, 21, 2, 22, 23, 24], simulating spectroscopies [25, 26, 27, 28, 29], and solving NP-complete problems [30, 31, 32, 33]. Among them, most of these applications use the variational technique to find a solution with a minimal value of the loss function or the energy, which is analogous to the strategy used in machine learning techniques, inspiring the idea of quantum machine learning [34, 35].
@@ I Introduction
Understanding the origin of superior performance in quantum machine learning is essential for optimizing quantum models and potentially enriching classical algorithms. Quantum-inspired classical algorithms, for example, have already accelerated linear algebra computations, showcasing the benefits of translating quantum insights to classical contexts [57, 58, 59, 60]. Despite these advancements, the machine learning field sees a paucity of quantum-inspired algorithms, primarily due to an inadequate grasp of quantum model intricacies [61]. This knowledge gap has spurred investigations into the expressibility of quantum models [62, 63, 64, 65]. However, uncovering the relationship between the expressibility and the architecture of a quantum circuit is challenging. Consequently, research has predominantly pursued numerical evaluations, including evaluating the expressibility of the Hilbert space and the functional expressibility [62, 63, 64, 65]. For instance, the expressibility of the Hilbert space was characterized by comparing different distributions of the fidelity between the quantum state from a quantum model and the Haar-random states [62]. In addition, Schude et al. proposed utilizing Fourier series to explain the functional expressibility of a quantum model [63]. While these insights advance our understanding, they have not elucidated the relationship between Hilbert space expressibility and quantum machine learning efficacy. Moreover, the decomposition of a function into a Fourier series cannot intuitively reflect the nonlinear relationship between the input data and the output. To advance quantum machine learning techniques, it is vital to address these foundational questions.
Our research advances the understanding of quantum models in quantum machine learning by providing clear analytical expressions for quantum circuits, distinguishing our work from previous studies. Specifically, we investigate two shallow quantum circuits integrated within a convolutional neural network (CNN) framework. Compared to classical CNNs, the hybrid quantum-classical CNN model requires significantly fewer training steps while achieving comparable training accuracy. Notably, we demonstrate that a special quantum activation function can efficiently extract edge information from an object while filtering out redundant data, resulting in faster convergence of the loss function. Additionally, we explore three quantum-inspired activation functions, all of which achieve high accuracy with fewer training steps compared to classical CNNs. These findings suggest that the higher-order nonlinear properties of quantum-inspired activation functions can enhance the performance of machine learning models. Furthermore, we develop a hybrid quantum Chebyshev-polynomial network (QCPN), capable of approximating any continuous function using only three neural network layers. In contrast, a standard three-layer classical neural network cannot achieve this level of approximation. This result highlights the potential of quantum-inspired activation functions to reduce the depth of machine learning models, offering a promising approach for optimizing large-scale models.
This paper is organized as follows. We highlight the major contributions and the importance of our work in Sec. II. In Sec.III, we present the functional formulation of a generic quantum circuit. In Sec.IV, we analyze the training results of a hybrid quantum-classical CNN using three distinct datasets. We also introduce three quantum-inspired activation functions and evaluate the performance of the CNN with these functions. Furthermore, we develop a quantum Chebyshev-polynomial network, capable of approximating any continuous function using three layers. Finally, in Sec. V, we discuss the future directions of quantum machine learning and provide perspectives on its ongoing development.
Figure 1: (a) Schematic of a generic quantum circuit, consisting of a data-encoding layer and a variational quantum layer. (b) Architecture of Quantum Circuit 1. (c) Architecture of Quantum Circuit 2. (d) One-qubit circuit used for data re-uploading.
II Contribution
The major contributions of our work include the following three aspects:
  • We explained the functional expression of a generic quantum circuit that can be integrated into the classical machine learning methods. This explanation is essential to our understanding of quantum machine learning techniques. Furthermore, our work opens a new research direction on exploring quantum-inspired classical activation functions.
  • We discovered an interesting quantum activation function, which can extract edge information and drop redundant information inside an object.
  • We developed a quantum Chebyshev-polynomial network, which can approximate any continuous function using three neural network layers. However, a classical three-layer neural network does not have this learnability. Therefore, our developed quantum Chebyshev-polynomial network could be adopted to reduce the complexity of deep learning models while maintaining the same learning capability.
III Functional Expressibility
To understand the mechanism of quantum neural networks, we analyze the functional expressibility of a generic quantum circuit defined in Fig. 1 (a). As a linear quantum model, we can divide it into two parts: data encoding and variational quantum layers. In the data encoding part, we transform real-world data ${\bf x}$ to a quantum state using quantum operator $U({\bf x})$, which consists of rotation gates. The variational quantum layer consists of $L$ rotation layers ($V_{l}({\bf\theta})$) and $L$ CNOT layers. By applying measurements with Pauli operators $M$, the output $O$ of this quantum circuit is
$O=\langle 0|U^{\dagger}({\bf x})V^{\dagger}({\bf\theta}_{1},{\bf\theta}_{2},\cdots,{\bf\theta}_{L})MV({\bf\theta}_{1},{\bf\theta}_{2},\cdots,{\bf\theta}_{L})U({\bf x})|0\rangle$ (1)
where $V({\bf\theta}_{1},{\bf\theta}_{2},\cdots,{\bf\theta}_{L})$ denotes the quantum operator in the variational circuit.
By further analyzing the matrix element of $U(\bf x)$, $V({\bf\theta}_{1},{\bf\theta}_{2},\cdots,{\bf\theta}_{L})$, and $M$, we find that $O$ can be rewritten as
$O=$ $\sum_{s}f_{s}({\bf\theta},{\bf x}),$ (2)
$f_{s}({\bf\theta},{\bf x})=$ $c_{s}\prod_{i\in\mathcal{L}_{s,1}}\mathrm{cos}(\theta_{i}/2+q_{i})\prod_{j\in\mathcal{L}_{s,2}}\mathrm{cos}(\theta_{j}/2+q_{j})\times$
$\prod_{k\in\mathcal{L}_{s,3}}\mathrm{cos}(\pi x_{k}/2+q_{k})\prod_{l\in\mathcal{L}_{s,4}}\mathrm{cos}(\pi x_{l}/2+q_{l}).$
Here, $\mathcal{L}_{\gamma}$ is a set of integers, and the value of $c_{s}$ is either 1 or -1. $q_{i}$ belongs to the group of {$20,\pi/$ 2}. Equation (2) implies that the output of a quantum circuit is a polynomial function whose component is a continued product of the cosine series.
To clarify this polynomial function, we investigate two quantum circuits as examples, which are plotted in Fig. 1 (b) and Fig. 1 (c). In the quantum circuit 1 (QC 1) shown in Fig. 1 (b), the data encoding is achieved via $R_{y}$ gates, and the variational quantum circuit is made of one $R_{y}$ rotation layer and one CNOT layer, in which the CNOT gate is applied between two adjacent qubits, which first across even-indexed qubits and then odd-indexed qubits. In the quantum circuit 2 (QC 2) shown in Fig. 1 (c), the CNOT layer is replaced by a layer that applies a CNOT gate between each two adjacent qubits. With some derivations (see details in the Appendix), it is found that the output of QC 1 can be written as
$O_{QC1}(\theta,x)=\left\{\begin{matrix}-\prod^{n}_{i\in\text{odd}}\text{cos}\alpha_{i}&n\in\text{odd}\\ -\text{cos}\alpha_{n}\prod_{i\in\text{odd}}^{n-3}\text{cos}\alpha_{i}&n\in\text{even}\end{matrix}\right.$ (4)
where $\alpha_{i}=\theta_{i}+\pi x_{i}$, and $n$ is the number of qubits. Additionally, the output of QC 2 can be expressed as
$O_{QC2}(\theta,x)=\left\{\begin{matrix}\left(\prod_{i=n-2}^{i=i-4,i>4}\text{cos}\alpha_{i}\right)\text{cos}\alpha_{3}\text{cos}\alpha_{1}&\text{when }n=4+4m\\ \left(\prod_{i=n-2}^{i=i-4,i>6}\text{cos}\alpha_{i}\right)\text{cos}\alpha_{4}\text{cos}\alpha_{2}&\text{when }n=6+4m\\ \left(\prod_{i=n-2}^{i=i-4,i>5}\text{cos}\alpha_{i}\right)\text{cos}\alpha_{5}\text{cos}\alpha_{2}&\text{when }n=3+4m\\ \left(\prod_{i=n-2}^{i=i-4,i>5}\text{cos}\alpha_{i}\right)\text{cos}\alpha_{3}\text{cos}\alpha_{1}&\text{when }n=5+4m\end{matrix}\right.$ (5)
where $m$ is a non-negative integer.
As observed from the equations above, the polynomial function transitions into a monotonic function for QC 1 and QC 2. In general, increasing the number of rotation layers in the variational quantum layers enhances the complexity of the output function by adding more components. This conclusion also applies to the data re-uploading technique illustrated in Fig. 1(d). In this method, the repeated application of rotation gates effectively increases the number of matrix multiplications, which, in turn, augments the number of components in the resulting polynomial function.
Figure 2: (a) Hybrid quantum-classical convolutional neural network (CNN) architecture. (b)-(d) Training loss curves for the CNN using various activation functions across epochs on the MNIST, FMNIST, and Letter datasets, respectively. (e)-(g) Comparison of the training and testing loss curves for the CNN using Activation Functions (AF) 1 and 3 across epochs for the MNIST, FMNIST, and Letter datasets, respectively. (h)-(j) Accuracy residuals for the CNN using AF 1 and AF 3 across epochs for both training and testing on the MNIST, FMNIST, and Letter datasets. (k) Summary table of the training accuracy, testing accuracy, and training steps for CNNs using different activation functions.
Figure 3: (a) The original input image and the corresponding feature maps in the first hidden layer of the convolutional neural network (CNN) using Activation Functions (AF) 1, 3, and 4, as shown in Fig. 2(k). (b)-(d) Trained angle parameters of the CNN with AF 3 for the MNIST, FMNIST, and Letter datasets, respectively. Different colors along the vertical direction denotes different features.
IV Quantum neural network
Quantum neural networks can be employed to perform various machine-learning tasks. However, their application to real-world problems is constrained by the limited availability of quantum resources. As a result, using purely quantum neural networks to solve problems involving high-dimensional data is currently impractical. To overcome this challenge, two primary strategies are often utilized. The first approach involves hybrid quantum-classical CNNs, which only require a small patch of an image to be processed by the quantum model. The second approach reduces the dimensionality of the input data using classical machine learning techniques such as principal component analysis (PCA) or autoencoders. In this situation, a pure quantum model can be adopted to solve the problem. However, we note that a simple classical two-layer neural network can be easily trained with high accuracy using these extracted features. It seems that adopting quantum models is not necessary for the second strategy. Therefore, our work will focus on the first strategy.
IV.1 Hybrid quantum-classical convolutional neural network